# Adaptive Transfer Clustering: A Unified Framework

### v2

@@ 2.1 Problem Setup
We could obtain the following expected misclassification rate for ˇ$\check{}\mathbf{Z}_{0}^{\textsf{ITL}}$ and ˇ$\check{}\mathbf{Z}_{0}^{\textsf{DP}}$, whose proof is deferred to Section C.2.
@@ 2.2 Transfer learning procedure
When $\varepsilon=0$, labels are perfectly matched and the error is purely caused by noise. In this case, the error bound given by Theorem 1 is $\psi\left(\lambda\right):=\mathcal{M}\left({\mu}/{\sigma},0,\lambda\right)$. When $\varepsilon>0$, the label discrepancy leads to an additional term $\phi\left(\lambda\right):=2\varepsilon\Phi\left(-\mu/\sigma+{\lambda}\left(2\mu/\sigma\right)^{-1}\right)$. Therefore, $\psi\left(\lambda\right)$ and $\phi\left(\lambda\right)$ can be interpreted as stochastic error and systematic error, respectively. Note that $\psi\left(\lambda\right)$ is non-increasing with $\lambda$ and $\phi\left(\lambda\right)$ is non-decreasing with $\lambda$, and the relation $\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\lambda\right)=\psi\left(\lambda\right)+\phi\left(\lambda\right)$ can be viewed as the bias-variance decomposition. In Theorem 3 below, we will show that choosing $\lambda=\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)$ as in (3) optimizes ${\cal M}\left(\mu/\sigma,\varepsilon,\lambda\right)$ up to a constant factor, and achieve the bias-variance trade-off. Formally, we have the following theorem regarding ˇ$\check{}\mathbf{Z}_{0}^{\lambda}$.
@@
We certify the optimality of our proposed transfer learning procedure (2.2) by establishing a matching lower bound for estimating $\mathbf{Z}_{0}^{*}$. The proof is deferred to Section B.3.
@@ 2.3 Adaptivity: Goldenshluger-Lepski method
We have the following theorem (which is a special case of Theorem 7) regarding the performance of ˇ$\check{}\mathbf{Z}_{0}^{\widehat{\lambda}}$, denoted by $\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\mu,\sigma\right)$. The proof can be found in Section B.4.
@@
for some universal constant $C_{1}>0$.
Due to the smoothness of $\Phi$, it follows that that $\check{\mathbf{Z}}_{0}^{\widehat{\lambda}}$ can achieve the optimal clustering error rate $\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\log\left({\varepsilon}^{-1}\left(1-\varepsilon\right)\right)\right)$, up to an additive factor of ${\log(M\zeta^{-1})}/n$, provided that the grid in $\Lambda$ is sufficiently dense. Formally, we have the following result regarding $\min_{\lambda\in\Lambda}\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\lambda\right)$, whose proof can be found in Section C.1.
@@
Suppose $M=\lceil\log^{2}n\rceil$ and $\lambda_{j}=jM^{-1}\log n$ for $j=0,\cdots,M$, we have
$\min_{\lambda\in\Lambda}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\lambda\right)\leq C_{0}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right),$
for some universal constant $C_{0}>0$.
@@ 3.1 Problem setup
$\log L_{\textsf{GMM}_{2}}\left(\mathbf{Z},\mathbf{X};\left(\mu,\sigma\right)\right)=\sum_{i=1}^{n}\left[\frac{(X_{0,i}-u\mu)^{2}}{2\sigma^{2}}+\log\left(1/2\right)\right].$
@@
$\log L_{\textsf{GMM}_{K}}\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\left\{{\bm{\mu}}_{k}\right\}_{k=1}^{K},\left\{\mathbf{\Sigma}_{k}\right\}_{k=1}^{K}\right)\right)=\sum_{i=1}^{n}\left[\log{\cal N}\left(\mathbf{X}_{i}\mid{\bm{\mu}}_{Z_{i}},\mathbf{\Sigma}_{Z_{i}}\right)+\log\pi_{Z_{i}}\right],$
@@
$\log L_{\textsf{LCM}}\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{P}\right)\right)=\sum_{i=1}^{n}\sum_{j=1}^{d}\left[X_{ij}\log P_{jZ_{i}}+\left(1-X_{ij}\right)\log\left(1-P_{jZ_{i}}\right)\right]+\sum_{i=1}^{n}\log\pi_{Z_{i}}.$
@@
$\log L_{\textsf{SBM}}\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B}\right)\right)=\sum_{1\leq i<j\leq n}\left[X_{ij}\log B_{Z_{i}Z_{j}}+\left(1-X_{ij}\right)\log\left(1-B_{Z_{i}Z_{j}}\right)\right]+\sum_{i=1}^{n}\log\pi_{Z_{i}}.$
@@ 3.2 General transfer clustering methodology
We extend ATC to fitting the general framework. For any nuisance parameter (^$(\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1})$, our transfer learning estimators for the clustering labels are defined as follows:
@@
Given a grid set $\Lambda\subset[0,\infty)$ and $\widehat{\psi}$ on $\Lambda$, we generalize the Goldenshluger-Lepski procedure in Section 2 to Algorithm 2, in order to select the optimal $\lambda$. Here $\widehat{\psi}$ is estimated by using a bootstrap procedure detailed in Algorithm 3, which can be regarded as a generalization to Algorithm 1. The final estimator is denoted by ^$\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}=\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1}\right)$.
@@ 4 General theory
4.1 Assumptions and general analysis
We assume $\left(\mathbf{Z}^{*}_{0},\mathbf{Z}^{*}_{1}\right)\sim\Pi$ where $\Pi$ satisfies the following conditions:
  • ${Z^{*}_{0,i}}\overset{i.i.d.}{\sim}\Pi_{0}$ and ${Z^{*}_{1,i}}\overset{i.i.d.}{\sim}\Pi_{1}$ for $i\in[n]$.
@@ 4.1 Assumptions and general analysis
$\log L_{\textsf{SBM},i}\left(Z_{i},X_{i};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B},\mathbf{Z}_{-i},\mathbf{X}_{-i}\right)\right)=\sum_{j>i}\left[X_{ij}\log B_{Z_{i}Z_{j}}+\left(1-X_{ij}\right)\log\left(1-B_{Z_{i}Z_{j}}\right)\right]+\log\pi_{Z_{i}}.$
@@
Note that the variance term ${\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ only involves a “clean” model (i.e., with perfectly matched cluster labels) and can be estimated by parametric bootstrap. Define (ˇ$\left(\check{}\mathbf{Z}_{0}^{\lambda},\check{}\mathbf{Z}_{1}^{\lambda}\right):=\textsf{TC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\lambda,\mathbf{\Theta}^{*}_{0},\mathbf{\Theta}^{*}_{1}\right)$. It can be readily seen that the bias term ${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)$ can be decomposed as
@@
Define $\xi_{\widehat{\psi}}:=\max_{\lambda\in\Lambda}\left[{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)-\widehat{\psi}\left(\lambda\right)\right]_{+}$, we have the following theorem, whose proof is given in Appendix D.2.
Lemma 2 (Oracle inequality for Goldenshluger-Lepski (GL)).
@@
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq 4\min_{\lambda\in\Lambda}\left\{\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right\}+3\xi_{\widehat{\psi}}$
@@
We cast the following assumption, which can be regarded as the general version of Assumption 4.
@@
The following lemma indicates the validity of using $\widehat{\psi}$ in Algorithm 3 as our estimator for $\psi$, whose proof is given in Section D.3.
@@
In Theorem 5, $\phi\left(\lambda\right)$ and $\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ can be viewed as the bias and the variance terms, respectively. The latter acts as an proxy for $\psi\left(\lambda\right)$. The additional term $\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)$ accounts for the error introduced by the parameter estimation, which, together with $n^{-1}\log\left(1/\zeta\right)$, is typically negligible. In light of this, our adaptive method achieves the bias-variance trade-off over the grid $\Lambda$.
@@ 4.2 Two-component symmetric $d$-dimensional GMM
We now illustrate our general result using two-component multivariate GMM. Similar to the univariate case, let $\{Z_{0,i}^{*}\}_{i=1}^{n}$ and $\{Z_{1,i}^{*}\}_{i=1}^{n}$ be i.i.d. Rademacher latent variables such that $\mathbb{P}\big{(}Z_{0,i}^{*}\neq Z_{0,i}^{*}\big{)}\leq\varepsilon$. In addition, we have $\mathbf{X}_{0,i}|Z_{0,i}^{*}\sim N_{d}(Z_{0,i}^{*}{\bm{\mu}},\sigma^{2}\mathbf{I}_{d})$ and $\mathbf{X}_{1,i}|Z_{1,i}^{*}\sim N_{d}(Z_{0,i}^{*}{\bm{\mu}},\sigma^{2}\mathbf{I}_{d})$ for some ${\bm{\mu}}\in\mathbb{R}^{d}$ and $\sigma>0$, where $d\geq 1$. We investigate the performance of our method for this multivariate GMM with unknown $\left({\bm{\mu}},\sigma,\varepsilon\right)$ in the regime $\left\|{\bm{\mu}}\right\|/\sigma\rightarrow\infty$ and $\left\|{\bm{\mu}}\right\|/\sigma<\sqrt{2\log n}$. Suppose we have any estimator $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$ that satisfies the following assumption:
@@
We can construct projected samples $\big{\{}\widetilde{X}_{0,i},\widetilde{X}_{1,i}\big{\}}_{i=1}^{n}$ using the estimated parameters $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$:
$\widetilde{X}_{0,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{0,i}}{\left\|\widehat{\bm{\mu}}^{\top}\right\|},\quad\widetilde{X}_{1,i}:=\frac{\widehat{\bm{\mu}}^{\top}\mathbf{X}_{1,i}}{\left\|\widehat{\bm{\mu}}^{\top}\right\|},\quad\forall i\in[n].$
Due to independence, we have
$\widetilde{X}_{0,i}|Z_{0,i}^{*}\sim N\big{(}Z_{0,i}^{*}\left\|\widehat{\bm{\mu}}\right\|^{-1}{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}},\sigma^{2}\big{)},\quad\widetilde{X}_{1,i}|Z_{1,i}^{*}\sim N\big{(}Z_{1,i}^{*}\left\|\widehat{\bm{\mu}}\right\|^{-1}{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}},\sigma^{2}\big{)},\quad\forall i\in[n].$
Denote $\widetilde{\mu}:=\left\|\widehat{\bm{\mu}}\right\|^{-1}{\widehat{\bm{\mu}}^{\top}{\bm{\mu}}}$, then the problem is reduced to the univariate case with samples $\big{\{}\widetilde{X}_{0,i},\widetilde{X}_{1,i}\big{\}}_{i=1}^{n}$ and unknown $\left(\widetilde{\mu},\sigma,\varepsilon\right)$. Note that we can view $\left(\left\|\widehat{\bm{\mu}}\right\|,\widehat{\sigma}\right)$ as fixed parameters, which serves as a good approximation of $\left(\left\|{\bm{\mu}}\right\|,\sigma\right)$ by 3. We thus have the following result regarding ^$\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}$ given by Algorithm 2 and Algorithm 3, using the projected samples (~$\big{(}\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1}\big{)}$ and estimated paramter $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$, i.e., ^$\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}=\textsf{ATC}\left(\widetilde{}\mathbf{X}_{0},\widetilde{}\mathbf{X}_{1};\Lambda,\left\|\widehat{\bm{\mu}}\right\|,\sigma\right)$.
@@
Suppose Assumption 3 holds with $\delta_{n}=o\left(\frac{1}{\left(\left\|{\bm{\mu}}\right\|/\sigma\right)^{3}}\wedge\left(\frac{\left\|{\bm{\mu}}\right\|/\sigma}{\log n}\right)^{2}\right)$. Let $\zeta\in(0,1)$, and assume that $B\geq C_{0}\zeta^{-2}\log\left(M\zeta^{-1}\right)$ for some universal constant $C_{0}>0$, then with probability at least $1-\zeta$, we have
@@
for some universal constant $C_{1}>0$.

### v3

@@ 2.1 Problem Setup
We could obtain the following expected misclassification rate for ˇ$\check{}\mathbf{Z}_{0}^{\textsf{ITL}}$ and ˇ$\check{}\mathbf{Z}_{0}^{\textsf{DP}}$, whose proof is deferred to Appendix C.2.
@@ 2.2 Transfer learning procedure
When $\varepsilon=0$, labels are perfectly matched and the error is purely caused by noise. In this case, the error bound given by Theorem 1 is $\psi\left(\lambda\right):=\mathcal{M}\left({\mu}/{\sigma},0,\lambda\right)$. When $\varepsilon>0$, the label discrepancy leads to an additional term $\phi\left(\lambda\right):=2\varepsilon\Phi\left(-\mu/\sigma+{\lambda}\left(2\mu/\sigma\right)^{-1}\right)$. Therefore, $\psi\left(\lambda\right)$ and $\phi\left(\lambda\right)$ can be interpreted as stochastic error and systematic error, respectively. Note that $\psi\left(\lambda\right)$ is non-increasing with $\lambda$ and $\phi\left(\lambda\right)$ is non-decreasing with $\lambda$, and the relation $\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\lambda\right)=\psi\left(\lambda\right)+\phi\left(\lambda\right)$ can be viewed as the bias-variance decomposition. In Theorem 3 below, we will show that choosing $\lambda=\log\left(\varepsilon^{-1}\left(1-\varepsilon\right)\right)$ as in (3) optimizes ${\cal M}\left(\mu/\sigma,\varepsilon,\lambda\right)$ up to a constant factor, and achieve the bias-variance trade-off. Formally, we have the following theorem regarding ˇ$\check{}\mathbf{Z}_{0}^{\lambda}$, whose proof can be found in Appendix B.2.
@@
We certify the optimality of our proposed transfer learning procedure (2.2) by establishing a matching lower bound for estimating $\mathbf{Z}_{0}^{*}$. The proof is deferred to Appendix B.3.
@@ 2.3 Adaptivity: Goldenshluger-Lepski method
This leads to our adaptive estimator
$\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\mu,\sigma\right)=\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},$
where ATC is a shorthand for Adaptive Transfer Clustering.
@@
The following theorem provides an error bound for ATC, whose proof can be found in Appendix B.4.
@@
for some universal constant $C_{1}>0$ determined by $C_{0}$.
Due to the smoothness of $\Phi$, it follows that that ATC can achieve the optimal clustering error rate $\mathcal{M}\left({\mu}/{\sigma},\varepsilon,\log\left({\varepsilon}^{-1}\left(1-\varepsilon\right)\right)\right)$, up to a negligible additive term ${\log(M\zeta^{-1})}/n$, provided that the grid in $\Lambda$ is sufficiently dense. Formally, we have the following result, whose proof can be found in Appendix C.1.
@@
Suppose the conditions of Theorem 4 hold. Let $M=\lceil\log^{2}n\rceil$ and $\lambda_{j}=jM^{-1}\log n$ for $j=0,\cdots,M$, then there exists some universal constant $C_{2}>0$ determined by $C_{1}$ such that with probability at least $1-\zeta$,
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{2}\left[\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right)+\frac{\log\left(M\zeta^{-1}\right)}{n}\right].$
Moreover, if $\textsf{SNR}<\left(\frac{1}{2}-c\right)\log n$ for some constant $c\in(0,1)$, then there exists some universal constant $C_{3}>0$ determined by $C_{2}$ such that with probability at least $1-n^{-10}$,
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{3}\mathcal{M}\left(\frac{\mu}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right).$ (7)
Corollary 1 delivers that ATC estimator achieves the error rate that matches the lower bound in Theorem 3.
@@ 3.1 Problem setup
$\log L_{\textsf{GMM}_{2}}\left(\mathbf{Z},\mathbf{X};\left(\mu,\sigma\right)\right)=\sum_{i=1}^{n}\left[\frac{(X_{0,i}-\mu)^{2}}{2\sigma^{2}}+\log\left(1/2\right)\right].$
@@
$\log L_{\textsf{GMM}_{K}}\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\left\{{\bm{\mu}}_{k}\right\}_{k=1}^{K},\left\{\mathbf{\Sigma}_{k}\right\}_{k=1}^{K}\right)\right):=\sum_{i=1}^{n}\left[\log{\cal N}\left(\mathbf{X}_{i}\mid{\bm{\mu}}_{Z_{i}},\mathbf{\Sigma}_{Z_{i}}\right)+\log\pi_{Z_{i}}\right],$
@@
$\log L_{\textsf{LCM}}$ $\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{P}\right)\right)$
$:=\sum_{i=1}^{n}\sum_{j=1}^{d}\left[X_{ij}\log P_{jZ_{i}}+\left(1-X_{ij}\right)\log\left(1-P_{jZ_{i}}\right)\right]+\sum_{i=1}^{n}\log\pi_{Z_{i}}.$
@@
$\log L_{\textsf{SBM}}$ $\left(\mathbf{Z},\mathbf{X};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B}\right)\right)$
$:=\sum_{1\leq i<j\leq n}\left[X_{ij}\log B_{Z_{i}Z_{j}}+\left(1-X_{ij}\right)\log\left(1-B_{Z_{i}Z_{j}}\right)\right]+\sum_{i=1}^{n}\log\pi_{Z_{i}}.$
@@ 3.2 General transfer clustering methodology
We will develop an adaptive transfer clustering method for the general framework. For any nuisance parameter (^$(\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1})$, our transfer learning estimators for the clustering labels are defined as follows:
@@
Given a grid set $\Lambda\subset[0,\infty)$ and $\widehat{\psi}$ on $\Lambda$, we generalize the Goldenshluger-Lepski procedure in Section 2 to Algorithm 2, in order to select the optimal $\lambda$. Here $\widehat{\psi}$ is estimated by using a bootstrap procedure detailed in Algorithm 3, which can be regarded as a generalization to Algorithm 1. The final estimator is denoted by
$\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1}\right)=\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}}.$
@@ 4 General theory
In this section, we conduct a theoretical analysis of the adaptive estimator $\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\widehat{}\mathbf{\Theta}_{0},\widehat{}\mathbf{\Theta}_{1}\right)$ based on Algorithm 2 and Algorithm 3. We start with general results and then derive sharp adaptivity guarantees under the two-component Gaussian mixture model in $\mathbb{R}^{d}$.
4.1 Assumptions and general theories
We impose the following assumptions on $\left(\mathbf{Z}^{*}_{0},\mathbf{Z}^{*}_{1}\right)$:
  • ${Z^{*}_{0,i}}\overset{i.i.d.}{\sim}\Pi_{0}$ and ${Z^{*}_{1,i}}\overset{i.i.d.}{\sim}\Pi_{1}$ for all $i\in[n]$, where $\Pi_{0}$ and $\Pi_{1}$ are probability distributions over $[K]$;
@@ 4.1 Assumptions and general theories
$\log L_{\textsf{SBM},i}$ $\left(Z_{i},X_{i};\left(\left\{\pi_{k}\right\}_{k=1}^{K},\mathbf{B},\mathbf{Z}_{-i},\mathbf{X}_{-i}\right)\right)$
$:=\sum_{j>i}\left[X_{ij}\log B_{Z_{i}Z_{j}}+\left(1-X_{ij}\right)\log\left(1-B_{Z_{i}Z_{j}}\right)\right]+\log\pi_{Z_{i}}.$
@@
Note that the variance term ${\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ only involves a “clean” model (i.e., with perfectly matched cluster labels) and can be estimated by parametric bootstrap if $\mathbf{\Theta}^{*}_{0}$ and $\mathbf{\Theta}^{*}_{1}$ are known. Define (ˇ$\left(\check{}\mathbf{Z}_{0}^{\lambda},\check{}\mathbf{Z}_{1}^{\lambda}\right):=\textsf{TC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\lambda,\mathbf{\Theta}^{*}_{0},\mathbf{\Theta}^{*}_{1}\right)$. It can be readily seen that the bias term ${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\lambda}\right)$ can be decomposed as
@@
Ideally, for Algorithm 2, we want an oracle inequality of the form
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\lesssim\min_{\lambda\in\Lambda}\{\phi(\lambda)+\psi(\lambda)\}+\text{overhead}.$
We present the following theorem of the same flavor. See Section D.2 for the proof.
Lemma 2 (Oracle inequality for Goldenshluger-Lepski (GL) method).
@@
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq 4\min_{\lambda\in\Lambda}\left\{\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right\}+3\xi_{\widehat{\psi}},$
where
$\xi_{\widehat{\psi}}:=\max_{\lambda\in\Lambda}\left[{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)-\widehat{\psi}\left(\lambda\right)\right]_{+}.$
If ${\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq\widehat{\psi}\left(\lambda\right)$ for all $\lambda\in\Lambda$, then $\xi_{\widehat{\psi}}=0$ and
${\mathfrak{D}}\left(\widehat{}\mathbf{Z}_{0}^{\widehat{\lambda}},\bar{}\mathbf{Z}_{0}^{\infty}\right)\leq 4\min_{\lambda\in\Lambda}\left\{\phi\left(\lambda\right)+\widehat{\psi}\left(\lambda\right)\right\}.$ (9)
In general, we would hope that ${\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ is dominated by $\widehat{\psi}\left(\lambda\right)$ plus a small quantity, so that $\xi_{\widehat{\psi}}$ is small. This requires $\widehat{\psi}$ to be sufficiently large. Meanwhile, we do not want it to be too large because that would make the upper bound in (9) become vacuous. Fortunately, choosing $\widehat{\psi}$ by quantiles of bootstrap samples does a good job in reconciling the two competing forces.
@@
We cast the following assumption on the parameter estimation error.
@@
The following lemma indicates the validity of using $\widehat{\psi}$ in Algorithm 3 as our estimator for $\psi$, whose proof is given in Appendix D.3.
@@
In Theorem 5, $\phi\left(\lambda\right)$ and $\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}_{0}^{\lambda},\bar{}\mathbf{Z}_{0}^{\infty}\right)$ can be viewed as the bias and the variance terms, respectively. The latter acts as a proxy for $\psi\left(\lambda\right)$. The additional term $\gamma\mathbb{E}{\mathfrak{D}}\left(\bar{}\mathbf{Z}^{\infty}_{0},\mathbf{Z}_{0}^{*}\right)$ accounts for the error introduced by the parameter estimation, which, together with $n^{-1}\log\left(1/\zeta\right)$, is typically negligible. In light of this, our adaptive method achieves the bias-variance trade-off over the grid $\Lambda$.
@@ 4.2 Two-component symmetric $d$-dimensional GMM
We now illustrate our general result using two-component symmetric multivariate GMM. Similar to the univariate case, let $\{Z_{0,i}^{*}\}_{i=1}^{n}$ and $\{Z_{1,i}^{*}\}_{i=1}^{n}$ be i.i.d. Rademacher latent variables such that $\mathbb{P}\big{(}Z_{0,i}^{*}\neq Z_{0,i}^{*}\big{)}\leq\varepsilon$. In addition, we have $\mathbf{X}_{0,i}|Z_{0,i}^{*}\sim N_{d}(Z_{0,i}^{*}{\bm{\mu}},\sigma^{2}\mathbf{I}_{d})$ and $\mathbf{X}_{1,i}|Z_{1,i}^{*}\sim N_{d}(Z_{0,i}^{*}{\bm{\mu}},\sigma^{2}\mathbf{I}_{d})$ for some ${\bm{\mu}}\in\mathbb{R}^{d}$ and $\sigma>0$, where $d\geq 1$. We investigate the performance of our method for this multivariate GMM with unknown $\left({\bm{\mu}},\sigma,\varepsilon\right)$ in the regime $\left\|{\bm{\mu}}\right\|/\sigma\rightarrow\infty$ and $\left\|{\bm{\mu}}\right\|/\sigma<\sqrt{2\log n}$. Suppose we have any estimator $\left(\widehat{\bm{\mu}},\widehat{\sigma}\right)$ that satisfies the following assumption:
@@
Applying TC to the two-component symmetric multivariate GMM for any $\lambda>0$ leads to
$\left(\widehat{Z}_{0,i}^{\lambda},\widehat{Z}_{1,i}^{\lambda}\right)$ $=\operatorname*{arg\,min}_{u,v\in\{\pm 1\}}\left\{-\left\langle\widehat{\bm{\mu}},u\mathbf{X}_{0,i}+v\mathbf{X}_{1,i}\right\rangle+\lambda\widehat{\sigma}^{2}\mathbb{I}\left(u\neq v\right)\right\},\quad\forall i\in[n].$
For any grid set a grid set $\Lambda\subset[0,\infty)$, the following theorem provides an error bound for ^$\widehat{}\mathbf{Z}^{\widehat{\lambda}}_{0}=\textsf{ATC}\left(\mathbf{X}_{0},\mathbf{X}_{1};\Lambda,\widehat{\bm{\mu}},\widehat{\sigma}\right)$, whose proof can be found in Appendix B.5.
@@
Let $\zeta\in(0,1)$. Suppose Assumption 3 holds with $\delta_{n}\leq c_{0}\left(\log n\right)^{-3}$ and assume $B\geq C_{0}\zeta^{-2}\log\left(M\zeta^{-1}\right)$ for some universal constants $c_{0},C_{0}>0$. There exists some universal constant $C_{1}>0$ determined by $c_{0}$ and $C_{0}$ such that with probability at least $1-\zeta$,
@@
Analogous to Corollary 1, we have the following result certifying the adaptivity of ATC, whose proof is almost the same as that of Corollary 1 and hence omitted.
Corollary 2.
Suppose the conditions of Theorem 6 hold. Let $M=\lceil\log^{2}n\rceil$ and $\lambda_{j}=jM^{-1}\log n$ for $j=0,\cdots,M$, then there exists some universal constant $C_{2}>0$ determined by $C_{1}$ such that with probability at least $1-\zeta$,
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{2}\left[\mathcal{M}\left(\frac{\left\|{\bm{\mu}}\right\|}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right)+\frac{\log\left(M\zeta^{-1}\right)}{n}\right].$
Moreover, if ${\left\|{\bm{\mu}}\right\|}^{2}/\sigma^{2}<\left(1-c\right)\log n$ for some constant $c\in(0,1)$, then there exists some universal constant $C_{3}>0$ determined by $C_{2}$ such that with probability at least $1-n^{-10}$,
(ˇ$\ell\left(\check{}\mathbf{Z}_{0}^{\widehat{\lambda}},\mathbf{Z}_{0}^{*}\right)\leq C_{3}\mathcal{M}\left(\frac{\left\|{\bm{\mu}}\right\|}{\sigma},\varepsilon,\log\left(\frac{1-\varepsilon}{\varepsilon}\right)\right).$ (11)